# PoLAR: Polar-Decomposed Low-Rank Adapter Representation

### v1

@@ 1 Introduction
Large language models (LLMs) transformed the field of machine learning through their proficiency in understanding and generating text, as well as handling complex multimodal tasks. Typically, those models require a number of parameters that is on the scale of billions, rendering fine-tuning of large-scale models infeasible in hardware-constrained setups. Consequently, parameter-efficient fine-tuning methods have been developed to enable adaptation of LLMs with limited resources (Houlsby et al., 2019; Li and Liang, 2021; Lester et al., 2021; Hu et al., 2022). One notable work is that of Hu et al. (2022), which performs fine-tuning by learning an additive low-rank update parameterized as the product of two low-rank matrices. Such Low-Rank Adapters (LoRA) have been the subject of study in a large body of work. Various improvements, ranging from initialization strategies (Meng et al., 2024; Li et al., 2025) and custom learning rate settings for stable feature learning (Hayou et al., 2024b) to alternative parameterizations (Liu et al., 2024a) have been proposed ever since.
Recent work attempts to overcome the low-rank constraint imposed by LoRA while preserving its parameter-efficiency (Xia et al., 2024; Lialin et al., 2024; Zhao et al., 2024; Huang et al., 2025; Jiang et al., 2024). The underlying premise is that the low-rank nature of the adapter limits its expressiveness. However, this premise is at odds with recent theoretical results that LoRA can approximate any target transformer model reasonably well under mild assumptions (Zeng and Lee, 2024). Additionally, Kalajdzievski (2023) shows that raising the nominal rank does little to improve performance. Taken together, these findings suggest that the low-rank space offers sufficient expressiveness, but the classical low-rank adapter formulation struggles to fully utilize this potential.
@@
To address this pathology, we put forth PoLAR, a co-design of architecture and optimizer that mitigates the directional diversity collapse, as shown in Fig. 1(c). On the architecture side, PoLAR facilitates effective exploitation of the linear algebraic rank by expressing the low-rank update as the product of two column-orthogonal direction matrices and a $r\times r$ scale matrix. This parameterization can also be interpreted from the perspective of decoupling direction and magnitude, which arises naturally when the low-rank factors are lifted via the polar decomposition. On the optimizer side, we apply methods from Riemannian optimization (Boumal, 2023). Theoretically, we demonstrate that our co-design enables exponentially faster convergence than vanilla LoRA on a canonical problem. To further enable training at scale, we draw on recent advances in infeasible methods for optimization over Stiefel manifolds (Ablin and Peyré, 2022). The resulting method is tailored to GPU hardware, catalyzing a speedup of at least 3$\times$ over feasible methods. In a nutshell, our contribution is four-fold:
@@
[…]
@@ 1.1 Related Work
LoRA. Low-Rank Adapters (Hu et al., 2022) are now a standard approach for fine-tuning large language models. There are various efforts to further enhance LoRA, e.g., via adaptivity (Zhang et al., 2023), chaining (Lialin et al., 2024; Xia et al., 2024), low-bit training (Li et al., 2024b), modifications for long sequences (Chen et al., 2024), regularization (Li et al., 2024a), weight decomposition (Liu et al., 2024a), or sparsity (Nikdan et al., 2024). For this work, LoRA-variants, which change the parameterization of $\Delta\mathbf{W}$ are particularly relevant. DoRA (Liu et al., 2024a) reparameterizes the low-rank update according to weight normalization (Salimans and Kingma, 2016), learning direction and magnitude of the update separately. AdaLoRA (Zhang et al., 2023) allocates the rank of adapters adaptively according to layer importance. It learns an SVD-type update with a diagonal matrix and two orthogonal low-rank factors, which are regularized with a Frobenius norm penalty. VeRA and LoRA-XS propose novel parameterizations to further shrink the number of trainable parameters (Kopiczko et al., 2024; Bałazy et al., 2024). As these lightweight methods achieve performance on par with LoRA, they indirectly suggest that LoRA’s expressiveness remains underutilized.
Optimization on Manifolds. Riemannian optimization generalizes gradient methods to smooth manifolds (Absil et al., 2008; Bonnabel, 2013; Boumal, 2023) and has proven useful in scenarios such as matrix completion (Vandereycken, 2013) and eigenvalue problems (Edelman et al., 1998). Despite the theoretical appeal, such methods are rarely applied in large-scale settings, as expensive retraction operations are required to ensure the feasibility of iterates. Ablin and Peyré (2022) propose an infeasible method for optimization on the orthogonal manifold that does not require retractions. Rather than ensuring the feasibility of all iterates, the specifically designed landing field guarantees convergence to the manifold in the limit. Gao et al. (2022) extend their analysis to the Stiefel manifold, and Schechtman et al. (2023) further generalize this method.
Confined by space, other closely related topics, such as discussions on theoretical works on LoRA prototype problems, can be found in Appendix B.
@@ 2.1 Overcoming Low Stable Rank with PoLAR
Figure 2: $\mathsf{sr}(\Delta\mathbf{W})$ of Llama-2-7B low-rank updates fine-tuned on commonsense reasoning tasks with rank 32.
@@
The directional diversity of an update matrix is measured by the average pairwise Euclidean distance of neurons when projected to the unit sphere. Note that this distance is within $[0,2]$ with the lower and upper bounds attained by a pair of collinear neurons, pointing in the same or opposite directions, respectively. Consequently, a distribution of pairwise distances with most mass at the ends of the interval can be interpreted as evidence for low directional diversity. As observed in Fig. 1(b), the LoRA update closely aligns along a single direction, with most neural updates being nearly collinear. Using the compact SVD $\Delta\mathbf{W}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$, it is possible to explain this observation and identify a low stable rank as a driver behind the lack of directional diversity. Here, $\mathbf{U}=[\mathbf{u}_{1},\dots,\mathbf{u}_{r}]\in\mathbb{R}^{m\times r}$ and $\mathbf{V}=[\mathbf{v}_{1},\dots,\mathbf{v}_{r}]\in\mathbb{R}^{n\times r}$ have orthonormal columns and $\mathbf{\Sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{r})\in\mathbb{R}^{r\times r}$ contains the singular values on its diagonal. Also, let $\mathbf{w}_{i}$ denote the $i$-th row of $\Delta\mathbf{W}$. With this notation, we can see that the direction of the LoRA update for the $i$-th neuron can be approximated by:
@@
where $(a)$ comes from $\mathsf{sr}(\Delta\mathbf{W})\approx 1$. Equation (1) suggests that the weight update of all neurons tends to strongly align with the direction of the leading right singular vector up to a sign flip, causing the two-cluster pattern in Fig. 1(b). Moreover, Fig. 2 demonstrates the wide applicability of this finding, as the LoRA updates across several datasets suffer from a very low stable rank. We provide more thorough empirical evidence of this collapse in directional diversity in Appendix F.1.
@@
To this end, we advocate to incorporate orthogonality directly into the architecture. This can be achieved with polar decomposition; see Appendix C.1 for a brief review. Applied to each low-rank factor, this decomposition yields $\mathbf{Z}_{1}=\mathbf{X}\mathbf{\Theta}_{1}$ and $\mathbf{Z}_{2}=\mathbf{Y}\mathbf{\Theta}_{2}$, separating them into PSD scale and column-orthogonal direction components with $\mathbf{X}\in\mathsf{St}(m,r)$, $\mathbf{Y}\in\mathsf{St}(n,r)$ and $\mathbf{\Theta}_{i}\in\mathbb{S}^{r}_{\succeq 0}$ for $i\in\{1,2\}$. The desirable orthogonality is thus naturally enforced through the manifold structure of $\mathbf{X}$ and $\mathbf{Y}$. Moreover, rather than relying on two individual scale matrices, it is more convenient to learn a joint $\mathbf{\Theta}\in\mathbb{R}^{r\times r}$ matrix for the overall update, which amounts to merging the product $\mathbf{\Theta}_{1}\mathbf{\Theta}_{2}^{\top}$. These considerations give rise to our Polar-decomposed Low-rank Adapter Representation (PoLAR):
@@
As a byproduct, PoLAR admits a natural interpretation in terms of a direction–magnitude decomposition. However, unlike alternative decompositions, such as the column-wise weight normalization used in DoRA (Salimans and Kingma, 2016; Liu et al., 2024a), PoLAR enforces orthogonality, substantially increasing the stable rank (see Fig. 2) and generating low-rank updates with more competitive downstream task performance (see later in Section 4).
@@ 2.2 Faster Optimization with PoLAR Parameterization
We now compare the convergence of PoLAR and LoRA on the problem of learning a low-rank adapter for a single layer with whitened data. As discussed in Arora et al. (2018); Zhang and Pilanci (2024); Li et al. (2025) and Appendix C.4, applying LoRA in this case is equivalent to matrix factorization under the asymmetric Burer-Monteiro (BM) parameterization (Burer and Monteiro, 2003)
@@
In problem (3), the matrix to be factorized (or the target matrix of LoRA; see Appendix C.4) is
@@
denoted as $\mathbf{A}\!\in\!\mathbb{R}^{m\times n}$, and $\mathbf{Z}_{1},\mathbf{Z}_{2}$ represent the LoRA weights. In light of the low-rank setting, $r\ll\min\{m,n\}$ is assumed. Problem (3) has been widely adopted as a testbed for developing advanced optimization schemes for LoRA; see, e.g., Zhang and Pilanci (2024) or Li et al. (2025).
@@ 3 Practical PoLAR for Scalable Fine-Tuning
To fine-tune the PoLAR parameterization of (2), we initialize $\mathbf{X}_{0}$ and $\mathbf{Y}_{0}$ randomly on the Stiefel manifold and set $\mathbf{\Theta}_{0}=\mathbf{0}$. At iteration $t$, we update $\mathbf{X}_{t}$ and $\mathbf{Y}_{t}$ with the landing field. Taking the update on $\mathbf{X}_{t}$ as an example, we use
@@ 4 Experiments
To demonstrate PoLAR’s effectiveness at scale, we fine-tune backbones spanning 350M–27B parameters on three benchmarks of varying difficulty. Our study includes both masked and autoregressive language models and evaluates their abilities in terms of general language understanding, commonsense reasoning, and multi-step mathematical problem solving. The code for our experiments is available at https://github.com/kcc-lion/polar/.
@@ 4.1 Commonsense Reasoning
Figure 4: Left two: Stable rank distribution of Llama-2-7B adapter layers trained with PoLAR and LoRA and different $r$ on the SIQA and HeSw datasets. See also Fig. 5. Right two: Stable rank dynamics of the fifth transformer block over the course of training by layer type. See also Fig. 7.
In Table 8 of Appendix F.4, our evaluation is further scaled up, where we report results for Gemma-3-12B (Gemma Team, 2025) on the commonsense reasoning benchmark using the same setup. PoLAR is observed to improve upon LoRA by more than a percentage point.
@@ 4.2 Mathematical Problem Solving
Table 4: Runtime ($\mu\mathrm{s}$) of landing (Land.) and retraction (Retr.) on a H100 with parameter size $(4096,r)$
@@ 4.4 Practical Runtime Improvement over Retraction
For $\tilde{\mathbf{X}}\in\mathbb{R}^{m\times r}$, a landing step requires $\mathcal{O}(m^{2}r)$ FLOPs while a retraction-based routine only requires $\mathcal{O}(mr^{2})$ operations. While landing offers no advantage in terms of FLOPs over retraction-based methods, the General Matrix Multiplications (GEMMs) used in a landing step achieve much higher FLOP utilization on GPU hardware than the SVD-based retraction routine. To quantify runtime differences, we benchmark the retraction and landing operations using the most frequently occurring attention and MLP weight dimensions in Llama-2-7B LoRA fine-tuning. As Table 4 shows, even for very small ranks of $r=4$, which make retraction relatively cheaper, usage of the landing method provides roughly a speedup of 3×3$\times$3 × with the advantage increasing to 18$\approx$ 18$\times$ 18 × for higher ranks. Appendix F.2 extends the study to a broader range of layer shapes and additionally compares memory and runtime overhead to LoRA and DoRA.
5 Conclusion
Low-rank adaptation has become the workhorse for efficient fine-tuning, yet our analysis revealed that the classical BM parameterization in LoRA often underutilizes its allocated subspace: the stable rank of the learned updates collapses, limiting expressive power. Building on this empirical finding, as well as theoretical insights from a canonical low-rank optimization problem, we introduced PoLAR, a reparameterization inspired by the polar decomposition that is coupled with a landing field optimization procedure. Across backbones from 350M to 27B parameters and tasks that span general language understanding, commonsense reasoning, and multi-step mathematics, PoLAR delivered consistent accuracy gains. The improvements often scale with adapter rank, validating that maintaining a higher stable rank translates into richer, more task-aligned updates. In practice, PoLAR’s reliance on nothing more than matrix multiplications implies that it maps cleanly onto GPU hardware, offering a drop-in replacement for existing LoRA pipelines.

### v2

@@ 1 Introduction
Large language models (LLMs) transformed the field of machine learning through their proficiency in understanding and generating text, as well as handling complex multimodal tasks. Typically, those models require a number of parameters that is on the scale of billions, rendering fine-tuning of large-scale models infeasible in hardware-constrained setups. Consequently, parameter-efficient fine-tuning methods have been developed to enable adaptation of LLMs with limited resources (Houlsby et al., 2019; Li and Liang, 2021; Lester et al., 2021; Hu et al., 2022). One notable work is that of Hu et al. (2022), which performs fine-tuning by learning an additive low-rank update parameterized as the product of two low-rank matrices. Such Low-Rank Adapters (LoRA) have been the subject of study in a large body of work. Various improvements, ranging from initialization strategies (Meng et al., 2024; Li et al., 2025a) and custom learning rate settings for stable feature learning (Hayou et al., 2024b) to alternative parameterizations (Liu et al., 2024a) have been proposed ever since.
Recent work attempts to overcome the low-rank constraint imposed by LoRA while preserving its parameter-efficiency (Xia et al., 2024; Lialin et al., 2024; Zhao et al., 2024; Huang et al., 2025; Jiang et al., 2024). The underlying premise is that the low-rank nature of the adapter limits its expressiveness. However, this premise is at odds with recent theoretical results that LoRA can approximate any target transformer model reasonably well under mild assumptions on the rank that depend on the relationship between the pre-trained and the target model (Zeng and Lee, 2024). Additionally, Kalajdzievski (2023) shows that raising the nominal rank does little to improve performance. Taken together, these findings suggest that the low-rank space offers sufficient expressiveness, but the classical low-rank adapter formulation struggles to fully utilize this potential.
@@
[…]
@@
Notation. Bold capital (lowercase) letters denote matrices (vectors); $(\cdot)^{\top}$ and $\|\cdot\|_{\mathsf{F}}$ refer to transpose and Frobenius norm of a matrix; $\|\cdot\|$ is the $\ell_{2}$ (spectral) norm of a vector (matrix); $\sigma_{i}(\cdot)$ and $\lambda_{i}(\cdot)$ denote the $i$-th largest singular value and eigenvalue, respectively. For a matrix $\mathbf{X}\in\mathbb{R}^{r\times r}$, let $\mathsf{Skew}(\mathbf{X})=\frac{1}{2}(\mathbf{X}-\mathbf{X}^{\top})$ be its skew-symmetric part. The set of matrices with orthonormal columns, i.e., the Stiefel manifold, is denoted as $\mathsf{St}(m,r):=\{\mathbf{X}\in\mathbb{R}^{m\times r}:\mathbf{X}^{\top}\mathbf{X}=\mathbf{I}_{r}\}$. The set of $r\times r$ positive semi-definite (PSD) matrices is denoted as $\mathbb{S}^{r}_{\succeq 0}:=\{\mathbf{X}\in\mathbb{R}^{r\times r}:\mathbf{X}\succeq 0\}$.
@@ 1.1 Related Work
Memory-Efficient Fine-Tuning. Low-Rank Adapters (Hu et al., 2022) are now a standard approach for fine-tuning large language models. There are various efforts to further enhance LoRA, e.g., via adaptivity (Zhang et al., 2023), chaining (Lialin et al., 2024; Xia et al., 2024), low-bit training (Dettmers et al., 2023; Li et al., 2024), modifications for long sequences (Chen et al., 2024), weight decomposition (Liu et al., 2024a), or sparsity (Nikdan et al., 2024). For this work, LoRA-variants that change the parameterization of $\Delta\mathbf{W}$ are particularly relevant. DoRA (Liu et al., 2024a) reparameterizes the low-rank update according to weight normalization (Salimans and Kingma, 2016), learning direction and magnitude of the update separately. AdaLoRA (Zhang et al., 2023) allocates the rank of adapters adaptively according to layer importance, learning an SVD-type update with a diagonal matrix and two orthogonal low-rank factors. A work concurrent to ours (Li et al., 2025b) proposes a reparameterization of LoRA that is similar to ours, using feasible methods from the literature of Riemannian optimization for training. VeRA and LoRA-XS propose novel parameterizations to further shrink the number of trainable parameters (Kopiczko et al., 2024; Bałazy et al., 2024). Building on top of the LoRA-XS parameterization, Ponkshe et al. (2024) develop a custom initialization scheme which is informed by the SVD of the gradient at the pre-trained model. Another closely related line of research is dedicated to the development of subspace optimizers which maintain the memory efficiency of low-rank adapters, while enabling full rank parameter updates. Such subspace methods project gradients to a low-rank subspace within which optimizer states are maintained (Zhao et al., 2024; Liang et al., 2024; Hao et al., 2024; Robert et al., 2025; Chen et al., 2025).
Optimization on Manifolds. Riemannian optimization generalizes gradient methods to smooth manifolds (Absil et al., 2008; Bonnabel, 2013; Boumal, 2023) and has proven useful in scenarios such as matrix completion (Vandereycken, 2013) and eigenvalue problems (Edelman et al., 1998). Despite the theoretical appeal, such methods are rarely applied in large-scale settings, as expensive retraction operations are required to ensure the feasibility of iterates. Ablin and Peyré (2022) propose an infeasible method for optimization on the orthogonal manifold that does not require retractions. Rather than ensuring the feasibility of all iterates, the specifically designed landing field guarantees convergence to the manifold in the limit. Gao et al. (2022) extend their analysis to the Stiefel manifold, and Schechtman et al. (2023) generalize the landing method to more general cases.
@@ 2.1 Overcoming Low Stable Rank with PoLAR
Figure 2: $\mathsf{sr}(\Delta\mathbf{W})$ of Llama-2-7B low-rank updates fine-tuned on commonsense tasks with rank 32.
@@
The directional diversity of an update matrix is measured by the average pairwise Euclidean distance of neurons when projected to the unit sphere. Note that this distance is within $[0,2]$ with the lower and upper bounds attained by a pair of collinear neurons, pointing in the same or opposite directions, respectively. Consequently, a distribution of pairwise distances with most mass at the ends of the interval can be interpreted as evidence for low directional diversity.
As observed in Fig. 1(b), the LoRA update closely aligns along a single direction, with most neural updates being nearly collinear. Using the compact SVD $\Delta\mathbf{W}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$, it is possible to explain this observation and identify a low stable rank as a driver behind the lack of directional diversity. Here, $\mathbf{U}=[\mathbf{u}_{1},\dots,\mathbf{u}_{r}]\in\mathbb{R}^{m\times r}$ and $\mathbf{V}=[\mathbf{v}_{1},\dots,\mathbf{v}_{r}]\in\mathbb{R}^{n\times r}$ have orthonormal columns and $\mathbf{\Sigma}=\operatorname{diag}(\sigma_{1},\dots,\sigma_{r})\in\mathbb{R}^{r\times r}$ contains the singular values on its diagonal. Also, let $\mathbf{w}_{i}$ denote the $i$-th row of $\Delta\mathbf{W}$. With this notation, we can see that the direction of the LoRA update for the $i$-th neuron can be approximated by:
@@
where $(a)$ comes from $\mathsf{sr}(\Delta\mathbf{W})\approx 1$. Equation (1) suggests that the weight update of all neurons tends to strongly align with the direction of the leading right singular vector up to a sign flip, causing the two-cluster pattern in Fig. 1(b). Moreover, Fig. 2 demonstrates the wide applicability of this finding, as the LoRA updates across several datasets suffer from a very low stable rank.
We provide more thorough empirical evidence of this collapse in directional diversity in Appendix F.1.
@@
To this end, we advocate to incorporate orthogonality directly into the architecture. This can be done by applying the polar decomposition (Golub and Van Loan, 2013) to each of the low-rank factors.
Definition 1 (Polar Decomposition).
Let $\mathbf{Z}_{1}\in\mathbb{R}^{m\times r}$ with $m\geq r$ and $\mathbf{Z}_{2}\in\mathbb{R}^{n\times r}$ with $n\geq r$. It admits factorizations $\mathbf{Z}_{1}=\mathbf{X}\mathbf{\Theta}_{1}$ and $\mathbf{Z}_{2}=\mathbf{Y}\mathbf{\Theta}_{2}$, where $\mathbf{X}\in\mathsf{St}(m,r)$  [^1] The Stiefel manifold is defined as $\mathsf{St}(m,r):=\{\mathbf{X}\in\mathbb{R}^{m\times r}:\mathbf{X}^{\top}\mathbf{X}=\mathbf{I}_{r}\}$. and $\mathbf{Y}\in\mathsf{St}(n,r)$ have orthonormal columns and $\mathbf{\Theta}_{1},\mathbf{\Theta}_{2}\in\mathbb{R}^{r\times r}$ are positive semi-definite, respectively.
When $r=1$, Definition 1 reduces to the familiar magnitude-direction decomposition of vectors.
The polar decomposition thus generalizes magnitude-direction decomposition to matrices, where the semi-orthogonal factor represents the directional component and the positive semi-definite factor captures the magnitude. Using this magnitude-direction decomposition, the desirable orthogonality is naturally enforced through the manifold structure of $\mathbf{X}$ and $\mathbf{Y}$. Moreover, rather than relying on two individual scale matrices, it is more convenient to learn a joint $\mathbf{\Theta}\in\mathbb{R}^{r\times r}$ matrix for the overall update, which amounts to merging the product $\mathbf{\Theta}_{1}\mathbf{\Theta}_{2}^{\top}$.
These considerations give rise to our Polar-decomposed Low-rank Adapter Representation (PoLAR):
@@
As a byproduct, PoLAR admits a natural interpretation in terms of a direction–magnitude decomposition. However, unlike alternative decompositions, such as the column-wise weight normalization used in DoRA (Salimans and Kingma, 2016; Liu et al., 2024a), PoLAR enforces orthogonality, substantially increasing the stable rank and generating low-rank updates with more competitive downstream task performance (see later in Section 4).
@@ 2.2 Faster Optimization with PoLAR Parameterization
We now compare the convergence of PoLAR and LoRA on the problem of learning a low-rank adapter for a single layer with whitened data. As discussed in Arora et al. (2018); Zhang and Pilanci (2024); Li et al. (2025a) and Appendix C.3, applying LoRA in this case is equivalent to matrix factorization under the asymmetric Burer-Monteiro (BM) parameterization (Burer and Monteiro, 2003)
@@
In problem (3), the matrix to be factorized (or the target matrix of LoRA; see Appendix C.3) is denoted as $\mathbf{A}\!\in\!\mathbb{R}^{m\times n}$, and $\mathbf{Z}_{1},\mathbf{Z}_{2}$ represent the LoRA weights. In light of the low-rank setting, $r\ll\min\{m,n\}$ is assumed. Problem (3) has been widely adopted as a testbed for developing advanced optimization schemes for LoRA; see, e.g., Zhang and Pilanci (2024) or Li et al. (2025a).
@@ 3 Practical PoLAR for Scalable Fine-Tuning
To fine-tune the PoLAR parameterization of (2), we initialize $\mathbf{X}_{0}$ and $\mathbf{Y}_{0}$ randomly on the Stiefel manifold and set $\mathbf{\Theta}_{0}=\mathbf{0}$. At iteration $t$, we update $\mathbf{X}_{t}$ and $\mathbf{Y}_{t}$ with the landing field. Taking the
@@
update on $\mathbf{X}_{t}$ as an example, we use
@@ 4 Experiments
To demonstrate PoLAR’s effectiveness at scale, we fine-tune backbones spanning 350M–27B parameters on three benchmarks of varying difficulty. Our study includes both masked and autoregressive language models and evaluates their abilities in terms of general language understanding, commonsense reasoning, and multi-step mathematical problem solving.
@@ 4.1 Commonsense Reasoning
Figure 4: Left two: Stable rank distribution of Llama-2-7B adapter layers trained with PoLAR and LoRA and different $r$ on the Social-IQA and HellaSwag datasets. See also Fig. 5. Right two: Stable rank dynamics of the fifth transformer block over the course of training by layer type. See also Fig. 7.
In Appendix F, we provide further experimental results for the commonsense reasoning benchmark, comparing PoLAR to GaLore (Zhao et al., 2024). In addition, we further scale up the experiment and report detailed results for Gemma-3-12B (Gemma Team, 2025), where PoLAR is observed to improve upon LoRA by more than a percentage point.
@@ 4.2 Mathematical Problem Solving
Table 4: Runtime ($\mu\mathrm{s}$) of landing (Land.) and retraction (Retr.) on a H100 GPU with parameter size $(4096,r)$
@@ 4.4 Practical Runtime Improvement over Retraction
For $\tilde{\mathbf{X}}\in\mathbb{R}^{m\times r}$, a landing step requires $\mathcal{O}(m^{2}r)$ FLOPs while a retraction-based routine only requires $\mathcal{O}(mr^{2})$ operations. While landing offers no advantage in terms of FLOPs over retraction-based methods, the General Matrix Multiplications (GEMMs) used in a landing step achieve much higher FLOP utilization on GPU hardware than the SVD-based retraction routine. To quantify runtime differences, we benchmark the retraction and landing operations using the most frequently occurring attention and MLP weight dimensions in Llama-2-7B LoRA fine-tuning. As Table 4 shows, even for very small ranks of $r=4$, which make retraction relatively cheaper, usage of the landing method provides roughly a speedup of $3\times$ with the advantage increasing to $\approx 18\times$ for higher ranks. Appendix F.2 extends the study to a broader range of layer shapes and additionally compares memory and runtime overhead to LoRA and DoRA.
5 Conclusion and Future Work
Low‑rank adaptation has become the workhorse for efficient fine‑tuning, yet our analysis revealed that the classical BM parameterization in LoRA often underutilizes its allocated subspace: the stable rank of the learned updates collapses, limiting expressive power. Building on this empirical finding, as well as theoretical insights from a canonical low‑rank optimization problem, we introduced PoLAR, a reparameterization inspired by the polar decomposition that is coupled with a landing field optimization procedure. Across backbones from 350M to 27B parameters and tasks that span general language understanding, commonsense reasoning, and multi‑step mathematics, PoLAR delivered consistent accuracy gains. The improvements often scale with adapter rank, validating that maintaining a higher stable rank translates into richer, more task‑aligned updates. In practice, PoLAR’s reliance on nothing more than matrix multiplications implies that it maps cleanly onto GPU hardware, offering a drop‑in replacement for existing LoRA pipelines.
Given the low stable rank of the LoRA update, a closer examination of the spectral dynamics of low-rank fine-tuning seems to be a promising direction for future work. While several works already study the spectral evolution of weights during ordinary training of deep neural networks (Boix-Adsera et al., 2023; Martin and Mahoney, 2021; Mousavi-Hosseini et al., 2023), the impact of the explicit low-rank constraint on the spectrum is not yet well understood. An investigation into the dynamics of the spectral distributions of PoLAR could bring more insights into the benefits of the parameterization.
Acknowledgements
We are grateful to Christopher Criscitiello for insightful discussions and thank the anonymous reviewers for helpful feedback. Kai Lion is supported by Swiss National Science Foundation (SNSF) Sinergia Funding No. 216600. Liang Zhang gratefully acknowledges funding by the Max Planck ETH Center for Learning Systems (CLS). Bingcong Li is supported by SNSF Project Funding No. 200021-207343. Niao He is supported by ETH research grant funded through ETH Zurich Foundations and SNSF Starting Grant. This work was supported as part of the Swiss AI Initiative by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a105 on Alps.